# Effective-medium approach to the resonance distribution of wave scattering in a random point field

### v2

@@ Appendix B Scattering by a sphere and resonances
where the generalized spherical Bessel functions are defined by
$h^{\pm(d)}_{\ell}(z)$ $=\Gamma(\tfrac{d}{2})\left(\frac{2}{z}\right)^{\frac{d-2}{2}}H^{\pm}_{\ell+\frac{d-2}{2}}(z)\>,$ (116)
$j^{(d)}_{\ell}(z)$ $=\frac{h^{+(d)}_{\ell}(z)+h^{-(d)}_{\ell}(z)}{2}\>,$
where $H^{\pm}_{\nu}(z)$ are the Hankel functions [55]. The function $j^{(d)}_{\ell}(z)$ is finite and regular at $z=0$, hence its presence in the inner region of the ansatz (115). Imposing the continuity condition on the wavefunction and its derivative at the boundary $r=R$, we quickly get the following expression for the scattering matrix element

### v3

@@ Appendix A Scattering by a sphere and resonances
where $\kappa=\sqrt{k^{2}-nF(k)}$ and the generalized spherical Bessel functions are defined by
$j^{(d)}_{\ell}(z)$ $=\Gamma(\tfrac{d}{2})\left(\frac{2}{z}\right)^{\frac{d-2}{2}}J_{\ell+\frac{d-2}{2}}(z)\>,$ (64)
$h^{\pm(d)}_{\ell}(z)$ $=\Gamma(\tfrac{d}{2})\left(\frac{2}{z}\right)^{\frac{d-2}{2}}H^{\pm}_{\ell+\frac{d-2}{2}}(z)\>,$
in terms of the standard Bessel function $J_{\nu}(z)$ and the standard Hankel functions $H^{\pm}_{\nu}(z)$ [55]. The function $j^{(d)}_{\ell}(z)$ is finite and regular at $z=0$, hence its presence in the inner region of the ansatz (63). Imposing the continuity condition on the wavefunction and its derivative at the boundary $r=R$, we get the following expression for the scattering matrix element